Vector Space closed under scalar multiplication but not under addition

I just need an example of a vector space that is closed under scalar multiplication but not under addition. That is all. Thanks for your wisdom.
Answer #1

Let {(x,y) in R2: y=2*x or y=3*x}. That's a pretty simple space: just two lines in R2.

It's closed under scalar multiplication: if (x,y) lies on one of these lines (take y=3x, w.l.o.g.), then (αx,αy) is on the same line, since y=3ximpliesαy=3αx.

However, it's not closed under addition. If (x1,y1) satisfies y=2x and(x2,y2) satisfies y=3x, then(x1+x2,y1+y2) generally does not satisfy either of these relationships, so it is outside the space. Take(x1=1,y1=2) and(x2=1,y2=3) as anexample:(x1+x2=2,y1+y2=5), and5≠2*2 and 5≠3*2.

answered by: cubanita